XJoltz:

Factor completely, then place the factors in the proper location on the grid. a 2 - 9a + 20 Yall better help

XJoltz:

This is karma on my side

SUBTONICate6:

lol i have answer

XJoltz:

ACTUALLY ANSWER

XJoltz:

WHAT IS IT

SUBTONICate6:

ok

SUBTONICate6:

la mordedura de 87

XJoltz:

OH *GASP* REALLY?!

SUBTONICate6:

@xjoltz wrote:
OH *GASP* REALLY?!
la mordedura de 87

XJoltz:

THE ANSWER IS "ok"

XJoltz:

WOWWWWWWWWWWWW SO SMART

XJoltz:

just leave

SUBTONICate6:

=(

SUBTONICate6:

El Soup

XJoltz:

bro

SUBTONICate6:

Banana Banana Banana Banana Banana

XJoltz:

actually answer

SUBTONICate6:

you dum

SUBTONICate6:

To factor the quadratic expression completely, we can use the factoring method or the quadratic formula. Let's use the factoring method: The quadratic expression is: a^2 - 9a + 20 To factor it, we need to find two numbers whose sum is -9 and whose product is 20. The numbers -4 and -5 satisfy these conditions because (-4) + (-5) = -9 and (-4) * (-5) = 20. Therefore, we can rewrite the expression as follows: a^2 - 9a + 20 = (a - 4)(a - 5) Now, let's place the factors in the proper location on the grid: | a - 4 | a - 5 |


| 4 - 4 | 4 - 5 | The factors (a - 4) and (a - 5) go in the corresponding cells on the grid.

XJoltz:

@subtonicate6 wrote:
Banana Banana Banana Banana Banana
HOW IS THAT RELEVANT, EVEN IN THE SLIGHTEST? HOW?

XJoltz:

@subtonicate6 wrote:
To factor the quadratic expression completely, we can use the factoring method or the quadratic formula. Let's use the factoring method: The quadratic expression is: a^2 - 9a + 20 To factor it, we need to find two numbers whose sum is -9 and whose product is 20. The numbers -4 and -5 satisfy these conditions because (-4) + (-5) = -9 and (-4) * (-5) = 20. Therefore, we can rewrite the expression as follows: a^2 - 9a + 20 = (a - 4)(a - 5) Now, let's place the factors in the proper location on the grid: | a - 4 | a - 5 | | 4 - 4 | 4 - 5 | The factors (a - 4) and (a - 5) go in the corresponding cells on the grid.
No ai stupid

SUBTONICate6:

@xjoltz wrote:
@subtonicate6 wrote:
Banana Banana Banana Banana Banana
HOW IS THAT RELEVANT, EVEN IN THE SLIGHTEST? HOW?
I GAVE ANSWER YOU ID I OT

SUBTONICate6:

@xjoltz wrote:
@subtonicate6 wrote:
To factor the quadratic expression completely, we can use the factoring method or the quadratic formula. Let's use the factoring method: The quadratic expression is: a^2 - 9a + 20 To factor it, we need to find two numbers whose sum is -9 and whose product is 20. The numbers -4 and -5 satisfy these conditions because (-4) + (-5) = -9 and (-4) * (-5) = 20. Therefore, we can rewrite the expression as follows: a^2 - 9a + 20 = (a - 4)(a - 5) Now, let's place the factors in the proper location on the grid: | a - 4 | a - 5 | | 4 - 4 | 4 - 5 | The factors (a - 4) and (a - 5) go in the corresponding cells on the grid.
No ai stupid
I wrote it =)

XJoltz:

@subtonicate6 wrote:
@xjoltz wrote:
@subtonicate6 wrote:
To factor the quadratic expression completely, we can use the factoring method or the quadratic formula. Let's use the factoring method: The quadratic expression is: a^2 - 9a + 20 To factor it, we need to find two numbers whose sum is -9 and whose product is 20. The numbers -4 and -5 satisfy these conditions because (-4) + (-5) = -9 and (-4) * (-5) = 20. Therefore, we can rewrite the expression as follows: a^2 - 9a + 20 = (a - 4)(a - 5) Now, let's place the factors in the proper location on the grid: | a - 4 | a - 5 | | 4 - 4 | 4 - 5 | The factors (a - 4) and (a - 5) go in the corresponding cells on the grid.
No ai stupid
I wrote it =)
st fu your stupid

SUBTONICate6:

mean

XJoltz:

great now I cursed

SUBTONICate6:

HE HE HEH HAW

XJoltz: